OpenStudy (anonymous):

(01.02) Evaluate the following expression using the values given. Find 2x2 - 2z4 + y2 - x2 + z4 if x = -4, y = 3, and z = 2. Numerical Answers Expected!

OpenStudy (anonymous):

2*-4*2-2*2*4+3*2--4*4

OpenStudy (anonymous):

-10 XD

OpenStudy (anonymous):

\[2\left( -4 \right)^{2}-2\left( 2 \right)^{4}+3^{2}-\left( -4 \right)^{2}+2^{4}\] =2*16-2*16+9+16+16 =32-32+9+16+16=41

OpenStudy (anonymous):

correction last two lines only negative sign change 32-32+9-16+16=9

OpenStudy (anonymous):

So the answer is 9?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

yw